# Approximation of multipartite quantum states: revised version with new applications

### v1

@@ 4 On continuity of the regularized relative entropy of $\pi$-entanglement
where $A_{\pi(j)}$ is the system obtained by joining those $A_{i}$ such that $i\in\pi(j)$, $\mathrm{conv}(\cdot)$ and $\mathrm{cl}(\cdot)$ denote, respectively, the convex hull and the closure w.r.t. the trace norm.
@@
Let $\rho$ be a state of a $n$-partite quantum system $A_{1}\ldots A_{n}$. For any partition $\pi$, the relative entropy of $\pi$-entanglement of $\rho$ is defined as
@@
It is shown in [14] that this infimum is always attained and that the state at which it is attained is unique provided that $\rho$ is faithful state.
If $\pi$ is the finest partition $\left\{\{1\},\ldots,\{n\}\right\}$ then $\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A_{1}...A_{n}})$ coincides with the convex set $\mathfrak{S}_{\rm sep}(\mathcal{H}_{A_{1}...A_{n}})$ of all (fully) separable states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{m}})$ and hence $E_{R,\pi}$ is the relative entropy of entanglement $E_{R}$ in this case.
The relative entropy of $\pi$-entanglement possesses basic properties of entanglement measures (convexity, selective LOCC-monotonicity, asymptotic continuity, etc.) and satisfies the inequality
$pE_{R,\pi}(\rho)+(1-p)E_{R,\pi}(\sigma)\leq E_{R,\pi}(p\rho+(1-p)\sigma)+h_{2}(p),$ (37)
@@
The relative entropy of $\pi$-entanglement is nonadditive on tensor products, its regularization is defined as
@@
where $\rho^{\otimes k}$ is treated as a state of the $n$-partite quantum system $A^{k}_{1}...A^{k}_{n}$, where $A^{k}_{j}$ denotes the system obtained by joining $k$ copies of $A_{j}$. The above limit exists due to Fekete’s lemma (cf.[6]), because the relative entropy of $\pi$-entanglement is subadditive on tensor products.
It is shown in [14] that the relative entropy of $\pi$-entanglement $E_{R,\pi}$ is lower semicontinuous function on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{m}})$ for any partition $\pi$. Then a criterion of local continuity of $E_{R,\pi}$ is obtained in [15], which allows us to derive several verifiable conditions of local continuity of $E_{R,\pi}$ for any partition $\pi$, in particular, to prove that continuity of $E_{R,\pi}$ on a subset of $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ follows from continuity of the quantum mutual information on this subset (easy verifiable conditions for local continuity of the quantum mutual information are presented in [33, Section V]).
The regularized relative entropy of $\pi$-entanglement $E_{R,\pi}^{\infty}$ belongs to the class $II_{n}^{n-1}$ (in terms of Section 2.3) for any partition $\pi$: the faithful $II_{n}^{n-1}$-type continuity bound for $E_{R,\pi}^{\infty}$ is presented in Proposition 39 in [15]. This allows us to apply the approximation technique considered in Section 3 to the function $E^{\infty}_{R,\pi}$.
The following theorem contains results about analytical properties the regularized relative entropy of $\pi$-entanglement $E^{\infty}_{R,\pi}$ (proved by using Theorem 3B).
Theorem 2.  Let
$\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})=\left\{\hskip 1.0pt\rho\in\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})\,|\,\rho_{A_{s}}\!\!\in\mathfrak{S}_{\rm\!\textsf{FA}\!}(\mathcal{H}_{A_{s}})\textrm{ for }\,n-1\,\textrm{ indexes}\hskip 1.0pts\hskip 1.0pt\right\}$ (40)
be the subset of $\,\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ consisting of states $\rho$ such that the FA-property holds for at least $\hskip 1.0ptn-1$ marginal states of $\rho$. Let $\pi$ be any partition of $\{1,2,..,n\}$.
A) The function $E^{\infty}_{R,\pi}$ is finite and lower semicontinuous on the set $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$. Moreover,
$\liminf_{k\to+\infty}E^{\infty}_{R}(\rho_{k})\geq E^{\infty}_{R}(\rho_{0})$ (41)
for arbitrary sequence $\{\rho_{k}\}\subset\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ converging to a state $\rho_{0}\in\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$.
B) Let $\{\rho_{k}\}$ be an arbitrary sequence of states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ converging to a state $\rho_{0}\in\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$. The limit relation
holds provided one of the following conditions is valid: [^14] The finiteness of $\,E_{R,\pi}(\rho_{0})$ and $\,I(A_{1}\!:...:\!A_{n})_{\rho_{0}}$ follows from the condition $\rho_{0}\in\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$.
  1. $\rm a)$
  2. $\rm b)$
    there exists
    $\lim_{k\to+\infty}E_{R,\pi}(\rho_{k})=E_{R,\pi}(\rho_{0})<+\infty;$ (44)
  3. $\rm c)$
    there exists
    $\lim_{k\to+\infty}I(A_{1}\!:...:\!A_{n})_{\rho_{k}}=I(A_{1}\!:...:\!A_{n})_{\rho_{0}}<+\infty.$ (45)
Remark 6. The conditions $\rm a)$,$\,\rm b)$ and $\rm c)$ in Theorem 4 are related as follows: Theorem 1 in [15] states that $\,\rm a)\Rightarrow b)$ in general and that $\,\rm a)\Leftrightarrow b)$ if the state $\rho_{0}$ is faithful; Proposition 2 in [15] states that $\,\rm c)\Rightarrow b)$.
Easy verifiable conditions for the validity of $\,\rm c)$ are presented in [30, Section 5].
Proof. A) Let $\rho_{0}$ be a state in $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$. Assume that the FA-property holds for the states $[\rho_{0}]_{A_{1}}$,…,$[\rho_{0}]_{A_{n-1}}$. The finiteness of $E^{\infty}_{R,\pi}(\rho_{0})$ follows from inequality (38), since the FA-property of a state implies finiteness of its entropy [32].
To prove relation (41) for $E^{\infty}_{R,\pi}$ we apply Theorem 3B with the family of maps $\{\widetilde{\Lambda}^{1,...,n-1}_{r}\}_{r}$ defined in (27). It is mentioned before that the function $E^{\infty}_{R,\pi}$ belongs to the class $II_{n}^{n-1}$. So, Theorem 3B along with Remark 3 imply that
$\lim_{r\to+\infty}E^{\infty}_{R,\pi}(\widetilde{\Lambda}^{1,...,n-1}_{r}(\rho_{0}))=E^{\infty}_{R,\pi}(\rho_{0}).$ (46)
Consider the function
$f_{r}(\rho)=E^{\infty}_{R,\pi}(\widetilde{\Lambda}^{1,...,n-1}_{r}(\rho))=E^{\infty}_{R,\pi}(\Phi^{1}_{r}\otimes...\otimes\Phi^{n-1}_{r}\otimes\mathrm{Id}_{A_{n}}(\rho)),$
on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$, where $\Phi^{s}_{r}$, $s=1,2,...,n-1$, are the channels defined after (27). Since all the output states of the channel $\Phi^{s}_{r}$ are supported by the finite-dimensional subspace $P^{s}_{r}(\mathcal{H}_{A_{s}})$, the function $f_{r}$ is continuous on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ for each $r$ (this follows from the continuity bound (151) for $E^{\infty}_{R,\pi}$ in [14, Section V] obtained by generalizing the Winter arguments from [41]). Hence, $f_{*}=\sup_{r}f_{r}$ is a lower semicontinuous function on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$.
Since $f_{*}$ does not exceed the function $E^{\infty}_{R,\pi}$ by the monotonicity of $E^{\infty}_{R,\pi}$ under local channels, we have
$f_{*}(\rho_{k})\leq E^{\infty}_{R,\pi}(\rho_{k})\quad\forall k\qquad\textrm{and}\qquad f_{*}(\rho_{0})=E^{\infty}_{R,\pi}(\rho_{0}),$
where the equality follows from (46). Thus, the lower semicontinuity of $f_{*}$ implies (41).
B) By Remark 4 it suffices to prove that condition $\rm b)$ implies (42). Let $p\in\mathbb{N}$ be arbitrary. Since the cone generated by the set $\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A^{p}_{1}...A^{p}_{m}})$ is a weak* closed subset of $\mathfrak{T}_{+}(\mathcal{H}_{A^{p}_{1}...A^{p}_{m}})$ for any partition $\pi$ [14] and the state $\omega^{\prime}\otimes\omega^{\prime\prime}$ lies in $\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A^{p}_{1}...A^{p}_{m}})$ for any states $\omega^{\prime}\in\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A^{t}_{1}...A^{t}_{m}})$ and $\omega^{\prime\prime}\in\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A^{p-t}_{1}...A^{p-t}_{m}})$, $t=1,2,...,p-1$, the condition $\rm b)$ and the repeated use Corollary 2G in [15] allow us to show that
It follows that
Remark 7. The proof of Theorem 4B shows that the limit relation (47) holds for an arbitrary sequence $\{\rho_{k}\}$ of states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ converging to any state $\rho_{0}$ in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ provided that one of the conditions $\rm a)$,$\,\rm b)$ and $\rm c)$ in Theorem 4B is valid.
5 Other applications
In this section we consider several results about properties of the relative entropy of $\pi$-entanglement and its regularization obtained by using Theorem 3B in Section 3.
5.1 Simplified definition of $E_{R,\pi}$
@@
The first result concerns the possibility to take the infimum in the definition (36) of the relative entropy of $\pi$-entanglement over a subset of $\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A_{1}...A_{n}})$ consisting of ”simple” $\pi$-separable states.
Proposition 3.  Let $A_{1},..,A_{n}$ be arbitrary quantum systems and $\,\pi=\{\pi(j)\}_{j=1}^{\ell}$ a given partition of the set $\{1,2,...,n\}$.
@@
If $\rho$ is a state from the set $\,\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$ defined in (40) then the infimum in the definition (36) of $E_{R,\pi}(\rho)$ can be taken only over all finitely-decomposable $\pi$-separable states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$, i.e. states having the form
$\sigma=\sum_{k=1}^{m}p_{k}\hskip 1.0pt|\psi^{k}_{1}\rangle\langle\psi^{k}_{1}|\otimes...\otimes|\psi^{k}_{\ell}\rangle\langle\psi^{k}_{\ell}|,\quad\psi^{k}_{j}\in\mathcal{H}_{A_{\pi}(j)},\;\;\langle\psi^{k}_{j}|\psi^{k}_{j}\rangle=1,\;\;m<+\infty,$ (48)
@@
where $\{p_{k}\}_{k=1}^{m}$ is a probability distribution and $A_{\pi(j)}$ is the system obtained by joining those $A_{i}$ such that $i\in\pi(j)$.
@@
Proposition 5.1 implies, in particular, that we may define $E_{R,\pi}(\rho)$ for any state $\rho$ in $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$ by ignoring the existence of countably-non-decomposable $\pi$-separable states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$, i.e. the states that can not represented as a countable convex mixture of pure $\pi$-separable states. This claim is not trivial, since in general there are no reasons to assume that the infimum in definition (36) of $E_{R,\pi}(\rho)$ can be taken over the dense subset of $\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A_{1}...A_{n}})$ consisting of countably decomposable $\pi$-separable states.
Proof. Let $\rho$ be a state in $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$. For each natural $s$ in $[1,n]$ let $P_{r}^{s}$ be the spectral projector of $\rho_{A_{s}}$ corresponding to its $r$ maximal eigenvalues (taking the multiplicity into account). It is mentioned in Section 4 that the function $E_{R,\pi}$ belongs to the class $II_{n}^{n-1}$. Hence Theorem 3B implies that
$\lim_{r\to+\infty}E_{R,\pi}(\rho_{r})=E_{R,\pi}(\rho),$ (49)
where $\rho_{r}=c_{r}^{-1}Q_{r}\rho\hskip 1.0ptQ_{r}$, $\,Q_{r}=P_{r}^{1}\otimes...\otimes P_{r}^{n}$, $c_{r}=\mathrm{Tr}Q_{r}\rho$.
Consider the function
@@
where $\widetilde{\mathfrak{S}}^{\pi}_{\mathrm{sep}}(\mathcal{H}_{A_{1}...A_{n}})$ is the set of finitely-decomposable separable states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$. Since the set $\widetilde{\mathfrak{S}}^{\pi}_{\mathrm{sep}}(\mathcal{H}_{A_{1}...A_{n}})$ is convex, the joint convexity of the relative entropy and Proposition 5.24 in [19] imply that the function $\widetilde{E}_{R,\pi}$ is convex and that the inequality (37) with $E_{R,\pi}$ replaced by $\widetilde{E}_{R,\pi}$ holds for any states $\rho$ and $\sigma$ in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$. Note also that
@@
where the second inequality follows from (4). It follows that the function $\widetilde{E}_{R,\pi}$ belongs to the class $L_{n}^{n-1}(2,1)\subset II_{n}^{n-1}$ (see Section 2.3). Hence for any state $\rho$ in $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$ Theorem 3B implies that
$\lim_{r\to+\infty}\widetilde{E}_{R,\pi}(\rho_{r})=\widetilde{E}_{R,\pi}(\rho).$
By Lemma 5.1 below, we may assume in the definition of $E_{R,\pi}(\rho_{r})$ that $\rho_{r}$ is a state of a finite-dimensional $n$-partite system. Hence, $\widetilde{E}_{R,\pi}(\rho_{r})=E_{R,\pi}(\rho_{r})$ for any $r$. So, the last limit relation and (49) show that $\widetilde{E}_{R,\pi}(\rho)=E_{R,\pi}(\rho)$ for any state $\rho$ in $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$. $\Box$
@@
Corollary 1. For an arbitrary $\pi$-separable state $\rho$ in $\,\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$ there is a sequence $\{\rho_{k}\}$ of $\pi$-separable states having the form (48) such that

### v2

@@ 4.1.1 Definitions and basic properties
where $A_{\pi_{k}(j)}$ is the system obtained by joining those $A_{i}$ such that $i\in\pi_{k}(j)$, $|\pi_{k}|$ – the number of elements in $\pi_{k}$, $\mathrm{conv}(\cdot)$ and $\mathrm{cl}(\cdot)$ denote, respectively, the convex hull and the closure w.r.t. the trace norm.
@@
The (unregularized) relative entropy of $\pi$-entanglement of a state $\rho$ of a $n$-partite system $A_{1}\ldots A_{n}$ corresponding to a (non-empty) set $\pi\subseteq P(n)$ is defined as
@@
It is shown in [6] that this infimum is always attained and that the state at which it is attained is unique provided that $\rho$ is a faithful state.
If $\pi$ consists of the single partition $\left\{\{1\},\ldots,\{n\}\right\}$ then $\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A_{1}...A_{n}})$ coincides with the convex set $\mathfrak{S}_{\rm sep}(\mathcal{H}_{A_{1}...A_{n}})$ of all (fully) separable states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ and hence $E_{R,\pi}$ is the relative entropy of entanglement $E_{R}$ in this case.
The relative entropy of $\pi$-entanglement possesses the basic properties of entanglement measures (convexity, selective LOCC-monotonicity, asymptotic continuity, etc.) and satisfies the inequality
$pE_{R,\pi}(\rho)+(1-p)E_{R,\pi}(\sigma)\leq E_{R,\pi}(p\rho+(1-p)\sigma)+h_{2}(p)$ (65)
@@
The relative entropy of $\pi$-entanglement is nonadditive on tensor products. If $\pi$ consists of a single partition then the relative entropy of $\pi$-entanglement is subadditive on tensor products, and hence one can define its regularization
@@
where $\rho^{\otimes k}$ is treated as a state of the $n$-partite quantum system $A^{k}_{1}...A^{k}_{n}$, where $A^{k}_{j}$ denotes the system obtained by joining $k$ copies of $A_{j}$. The above limit exists due to Fekete’s lemma [47].
For the results below it is essential that for any permutation $\{s(i)\}_{i=1}^{n}$ of $[n]$ the following claims are valid:
  • the relative entropy of $\pi$-entanglement $E_{R,\pi}$ belongs to the class $II_{n}^{n-1}$ in the settings $A^{\prime}_{1}=A_{s(1)}$,…,$A^{\prime}_{n}=A_{s(n)}$ for any set of partitions $\pi$ (see Section 2.3);
  • the regularized relative entropy of $\pi$-entanglement $E_{R,\pi}^{\infty}$ belongs to the class $II_{n}^{n-1}$ in the settings $A^{\prime}_{1}=A_{s(1)}$,…,$A^{\prime}_{n}=A_{s(n)}$ for any $\pi$ consisting of a single partition.
These claims follow from Proposition 39 in [6] (where a common faithful $II_{n}^{n-1}$-type continuity bound for $E_{R,\pi}$ and $E_{R,\pi}^{\infty}$ is presented) and Remark 4.1.1 below.
Remark 11. In general, the functions $E_{R,\pi}$ and $E_{R,\pi}^{\infty}$ are not symmetrical w.r.t. the subsystems $A_{1}$,…,$A_{n}$. But as long as we believe that $\pi$ is arbitrary, we can assume that the energy constraints are imposed on the first $n-1$ subsystems $A_{1}$,…,$A_{n-1}$. Indeed, any permutation $A^{\prime}_{1}=A_{s(1)}$,…,$A^{\prime}_{n}=A_{s(n)}$ of the subsystems can be taken into account by the appropriate replacement $\pi\rightarrow\pi^{\prime}$ of the set of partitions.
4.1.2 Simplified definition of $E_{R,\pi}$
In this subsection and below we apply the results from Section 3 for exploring properties of the functions $E_{R,\pi}$ and $E_{R,\pi}^{\infty}$ in the infinite-dimensional settings (extending the results obtained in [6, 7]).
@@ 4.1.2 Simplified definition of $E_{R,\pi}$
The following proposition concerns the possibility to take the infimum in the definition (64) of the relative entropy of $\pi$-entanglement over a subset of $\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A_{1}...A_{n}})$ consisting of ”simple” $\pi$-separable states.
Proposition 4.  Let $A_{1},..,A_{n}$ be arbitrary quantum systems and $\,\pi=\{\pi_{k}\}_{k\in K_{\pi}}$ a given set of partitions of $\{1,2,...,n\}$. Let
[…]
be the set consisting of states $\rho$ of the $n$-partite system $A_{1}...A_{n}$ such that the FA-property holds for at least $\hskip 1.0ptn-1$ marginal states of $\rho$.
@@
If $\rho$ is a state in $\,\widetilde{\mathfrak{S}}^{n-1}_{II}(\mathcal{H}_{A_{1}...A_{n}})$ then the infimum in the definition (64) of $E_{R,\pi}(\rho)$ can be taken only over all finitely-decomposable $\pi$-separable states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$, i.e. states having the form
$\sigma=\sum_{i=1}^{m}p_{i}\,\sigma^{i}_{1}\otimes\sigma^{i}_{2}\otimes...\otimes\sigma^{i}_{|\pi_{k(i)}|},\;\;m<+\infty,$ (69)
@@
where $\{p_{i}\}_{i=1}^{m}$ is a probability distribution, $k(i)\in K_{\pi}$ and $\sigma^{\hskip 1.0pti}_{j}$ is a state in $\mathcal{H}_{A_{\pi_{k(i)}(j)}}$ for all $\,i=\overline{1,m}\,$ and $\,j=\overline{1,|\pi_{k(i)}|}$ ($A_{\pi_{k(i)}(j)}\,$ is the system obtained by joining those $A_{l}$ such that $\,l\in\pi_{k(i)}(j)$).
@@
Proposition 4.1.2 implies, in particular, that we may define $E_{R,\pi}(\rho)$ for any state $\rho$ in $\,\widetilde{\mathfrak{S}}^{n-1}_{II}(\mathcal{H}_{A_{1}...A_{n}})$ by ignoring the existence of countably-non-decomposable $\pi$-separable states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$, i.e. the states that can not represented as a countable convex mixture of pure $\pi$-separable states. This claim is not trivial, since in general there are no reasons to assume that the infimum in definition (64) of $E_{R,\pi}(\rho)$ can be taken over the dense subset of $\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A_{1}...A_{n}})$ consisting of countably decomposable $\pi$-separable states (because the relative entropy is not a continuous function of its arguments).
Proof. By the arguments in Remark 4.1.1 we may assume that $\rho$ is a state from the set $\mathfrak{S}^{n-1}_{II}(\mathcal{H}_{A_{1}...A_{n}})$ defined in (40). Consider the function
@@
where $\widetilde{\mathfrak{S}}^{\pi}_{\mathrm{sep}}(\mathcal{H}_{A_{1}...A_{n}})$ is the subset of $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ consisting of all finitely-decomposable $\pi$-separable states. Since the set $\widetilde{\mathfrak{S}}^{\pi}_{\mathrm{sep}}(\mathcal{H}_{A_{1}...A_{n}})$ is convex, the joint convexity of the relative entropy and Proposition 5.24 in [48] imply that the function $\widetilde{E}_{R,\pi}$ is convex and that the inequality (65) with $E_{R,\pi}$ replaced by $\widetilde{E}_{R,\pi}$ holds for any states $\rho$ and $\sigma$ in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ and any $p\in[0,1]$. Note also that
@@
where the second inequality follows from (9). Hence, the function $\widetilde{E}_{R,\pi}$ belongs to the class $L_{n}^{n-1}(2,1)\subset II_{n}^{n-1}$ (see Section 2.3).
Thus, the equality
$E_{R,\pi}(\rho)=\widetilde{E}_{R,\pi}(\rho)$ (70)
for any state $\rho\in\mathfrak{S}^{n-1}_{II}(\mathcal{H}_{A_{1}...A_{n}})$ can be derived from claim $\rm(iv)$ of the $II_{n}^{m}$-part of Corollary 3.2 in Section 3.2 applied to the function $E=E_{R,\pi}\in L_{n}^{n-1}(1,1)\subset II_{n}^{n-1}$ by proving that (70) holds for any state $\rho\in\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ such that $\mathrm{rank}\rho_{A_{i}}<+\infty$ for all $i=1,2,...,n$. The validity of (70) for any such state $\rho$ follows from Lemma 4.1.2 below, since it shows that in the definition of $E_{R,\pi}(\rho)$ we may assume that $\rho$ is a state of a finite-dimensional $n$-partite quantum system. $\Box$
@@
Corollary 3. For an arbitrary $\pi$-separable state $\rho$ in $\,\widetilde{\mathfrak{S}}^{n-1}_{II}(\mathcal{H}_{A_{1}...A_{n}})$ there is a sequence $\{\rho_{k}\}$ of $\pi$-separable states having the form (69) such that